OpenStudy (anonymous):

simplify (4 + 3√3)(9 - 2√3)

OpenStudy (anonymous):

\[ 18+19 \sqrt{3} \]

jhonyy9 (jhonyy9):

36-8sqrt3+27sqrt3-18=18+19sqrt3